OpenStudy (anonymous):

HELP PLEASE @mathstudent55 Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of f[g(3)]? 39 −3 −7 −1

OpenStudy (johnweldon1993):

First...what is g(3)? *plug 3 in for every 'x' you see in the equation*

OpenStudy (anonymous):

Im not sure

OpenStudy (johnweldon1993):

g(x) = x² - 10 What you do *because you want g(3) is replace every 'x' you see with a 3...so g(3) = (3)² - 10 g(3) = 9 - 10 g(3) = ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

So d

OpenStudy (johnweldon1993):

Right -1 ...so now......What you do...is take that -1 ....and plug it in for 'x' in your f(x) equation f(x) = 2x + 1 f(-1) = 2(-1) + 1 f(-1) = -2 + 1 f(-1) = ???

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

That my answer

OpenStudy (johnweldon1993):

Not quite...remember -2 + 1....you are adding 1 to -2 so it becomes -1 right?

OpenStudy (anonymous):

-1 sorry

OpenStudy (anonymous):

Can you help with two more questions please?

OpenStudy (johnweldon1993):

No problem! Good work! So in general...with an f(g(number)) problem like that.... 1) solve for g(number) 2) take that answer and plug it into the 'f' equation for 'x' 3) solve that and you have your answer

OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of (f − g)(−2)? −8 −5 −11 3

OpenStudy (johnweldon1993):

Okay now you take each equation seperatly and solve it with (-2) in place of 'x' f(x) = 3x - 2 f(-2) = 3(-2) - 2 f(-2) = -6 - 2 f(-2) = ??? and g(x) = x² - 1 g(-2) = (-2)² - 1 g(-2) = 4 - 1 g(-2) = ??? Then ...once you solve those...you take what you got from 'f'...and what you got from 'g'....and subtract them...

OpenStudy (anonymous):

So 3

OpenStudy (anonymous):

One More please

OpenStudy (anonymous):

Choose the best description of how to calculate f[g(3)] if f(x) = 3x − 2 and g(x) = x2 − 1. Calculate f(3) and g(3), then multiply them together. Calculate f(3) and g(3), then write the results as an ordered pair. Calculate f(3) and g(3), then subtract them. Calculate g(3), plug the result into f(x) and simplify.

OpenStudy (johnweldon1993):

I don't get 3 for the first one.... and for this new one...what did we do above before???

OpenStudy (johnweldon1993):

What do you get for f(-2) and what do you get for g(-2) ??

OpenStudy (anonymous):

SO 3 isnt the anser

OpenStudy (anonymous):

From earlier

OpenStudy (johnweldon1993):

No...like I asked....what do you get for 'f' and for 'g' when you do f(x) = 3x - 2 f(-2) = 3(-2) - 2 f(-2) = -6 - 2 f(-2) = ??? and g(x) = x² - 1 g(-2) = (-2)² - 1 g(-2) = 4 - 1 g(-2) = ???

OpenStudy (anonymous):

Im still lost

OpenStudy (anonymous):

So 3